6. Two cars are traveling along a straight road. Car (A) maintains a constant speed of 80.0 km/h; car (B) maintains a constant speed of 115 km/h. At t = 0, car (B) is 31.0 km behind car A. How far will car (A) travel from t-0 before it is overtaken by car (B)? km 7. A ball is thrown vertically downward from the edge of a bridge 25.0 m high with an initial speed of 18.0 m/s. The ball falls all the way down and strikes the water below. Determine the time it takes the ball to strike the water. 8. A photographer in a helicopter ascending vertically at a constant rate of 1.3 m/s accidentally drops a camera out the window when the helicopter is 55.0 m above the ground. (a) How long will it take the camera to reach the ground? 9.(b) What will its speed be just before it hits the ground? m/s 10. An object initially at rest experiences an acceleration of 1.50 m/s2 for 6.50 s then travels at that constant velocity for another 8.10 s. What is the magnitude of the objects average velocity over the 14.6 s interval? m/s
Answer #1

6)

115 x t = 80 t + 31

115t - 80t = 31

35t = 31

t = 31/35 = 0.89 hr

So, 80 x 0.89 = 70.86 km

7)

Using 2nd equation of motion,

-25 = (18 x t) - (0.5 x 9.8 t^2)

4.9t^2 - 18t - 25 = 0

Time, t = 4.75 sec

8)

V^2 = u^2 + 2as

V^2 = 1.3^2 + 2 x 9.8 x 55

V = 32.86 m/s

9)

V = u + at

-32.86 = 1.3 - 9.8t

Time, t = 3.49 sec

10)

D1 = 0.5 x 1.5 x 6.5^2 = 31.69 m

V = at = 1.5 x 6.5 = 9.75 m/s

D2 = 9.75 x 8.1 = 78.98 m

Average speed = (78.98 + 31.69)/14.6 = 7.58 m/s

Comment in case any doubt please rate my answer.....